# Propagation of Precessing Jet in Envelope of Tidal Disruption Events

### v1

@@ 1 Introduction
A star would be disrupted by a supermassive black hole (SMBH) when it comes too close to the central SMBH. In such events, namely tidal disruption events (TDEs), the stellar debris are eventually accreted by the SMBH, producing bright multi-band emission (Hills, 1975; Rees, 1988). The initial stellar orbit is most likely misaligned with the equatorial plane of the spinning SMBH. The tilted disk formed after the TDE would undergo Lense-Thirring precssion around the SMBH spin axis (e.g., Papaloizou & Lin, 1995; Fragile et al., 2007; Stone & Loeb, 2012; Lei et al., 2013; Wang et al., 2014; Franchini et al., 2016). The $\sim 15$ day quasi-periodic modulations found in the high-cadence X-ray monitoring observations of TDE candidate AT2020ocn, suggest the Lense-Thirring precession of a newly formed accretion disk (Pasham et al., 2024).
@@
For jetted TDEs, a question emerges: why would only a small fraction of TDEs successfully launch jets? One factor is the narrow beaming cone of the jets, and our line of sight is usually away from such jet cones. A recent study on IGR J12580+0134 revealed the first TDE with an off-axis relativistic jet (Lei et al., 2016; Yuan et al., 2016). Dai et al. (2018) proposed a unified model based on 3D general relativistic radiation magneto-hydrodynamis (GRRMHD) simulation that can explain both the optical and X-ray TDEs. The observed emission type mainly depends on the viewing angle of the observer with respect to the disk orientation. X-ray emissions will be observed from the optically thin funnel surrounded by winds when viewed face-on. For edge-on observations, only optical/UV emissions can be detected due to the heavily reprocessing of X-ray emissions in the chick wind and disk (Dai et al., 2018; Thomsen et al., 2022). However, even after the correction of relativistic beaming of jet and view angle effects, the fraction of jetted TDEs ($10^{-3}$ to $10^{-2}$ of all TDEs) is still much smaller than the fraction of active galactic nuclei with jets ($\sim 10$ per cent) (Teboul & Metzger, 2023; Lu et al., 2024). This remains an open question for jetted TDEs.
@@ 2.1 ZEBRA Envelope
where q=3/2nq=3/2-n, $\alpha=\frac{1-q(\gamma-1)}{\gamma-1}$, $\beta=\frac{\gamma-1}{1+\gamma-q(\gamma-1)}$, and $a=\frac{2-2q(\gamma-1)}{1+\gamma-q(\gamma-1)}$. $r_{0}\sim\chi\frac{2GM_{\bullet}}{c^{2}}$ is the inner radius of the disk, and $\chi$ is order of a few. $\rho_{0}$ is the density at this radius. The parameter $n$ is defined by Blandford and Begelman so that the accretion rate $\propto r^{n}$.
@@
Fig. 1 illustrates the time evolution of the $q$ and the global properties of the ZEBRA envelope. We consider the disruption of star with mass $M_{*}=M_{\odot}$ (solid lines) and $M_{*}=0.5M_{\odot}$ (dashed lines) by SMBH with mass $M_{\bullet}=10^{5}M_{\odot}$ (black lines) and $M_{\bullet}=10^{6}M_{\odot}$ (red lines). We show the evolution of $q$ in the panel (a). One can find that $M_{\bullet}$ and $M_{*}$ have significant effects for the whole evolution of $q$. In the panels (b)-(e), we also display the evolution of $T_{\rm env}$, $\rho_{0}$, $L_{\rm j}$ and $r_{\rm out}$ for different $M_{*}$ and $M_{\bullet}$.
@@ 2.2 Precessing Jet
Figure 2: A schematic picture of the precession jet in an inclined ZEBRA envelope. The jet precess with angle $\theta_{\rm LT}$. The angle between the SMBH spin axis ($\Omega_{\bullet}$) and the envelope’s rotation axis ($\Omega_{\rm env}$), i.e., the envelope inclination, is $\theta_{\rm env}$. Left panel: $\theta_{\rm LT}=\theta_{\rm env}$. In the low density polar regions, i.e, “free zone”, the relativistic jet can escape freely when it precess to this direction. The jet is choked at other regions. Middle and right panels: $\theta_{\rm LT}<\theta_{\rm env}$. The jet will be choked in all directions if $\theta_{\rm LT}$ is large (middle panel), or break out of envelope if $\theta_{\rm LT}$ is small (left panel).
Now, we introduce a precessing jet (with opening angle $\theta_{\rm j}$) launching from the center of the inclined ZEBRA envelope. The inner accretion disk undergoes Lense-Thirring precession with an precessing angle $\theta_{\rm LT}\leqslant\theta_{\rm env}$. We assume that the jet is normal to the inner disk and precess with the same period $P_{\rm LT}$. We show the schematic picture of our model in Fig. 2.
@@
The jet duration is $T_{\rm j}=\xi_{\rm duty}P_{\rm LT}$ and the width of the shell is $\Delta_{\rm j}\approx T_{\rm j}c$. The jet would be choked if its tail catches up to the head insider the ZEBRA envelope photosphere $r_{\rm ph}$ (defined by $\int_{r_{\rm ph}}^{r_{\rm out}}\kappa\rho dr=1$). Assuming the catch-up time $t_{\rm c}$ in the lab frame and the catch-up radius $r_{\rm c}$, one finds $(v_{\rm t}-v_{\rm h})t_{\rm c}=\Delta_{\rm j}$, where $v_{\rm t}$ and $v_{\rm h}$ are the tail and head velocity of the jet, and $\Delta_{\rm j}\approx cT_{\rm j}$ is the length of the jet. The jet catch-up time is then $t_{\rm c}=\Delta_{\rm j}/(v_{\rm t}-v_{\rm h})\simeq 2\Gamma_{\rm j}^{2}T_{\rm j}$, and $\Gamma_{\rm j}$ is the jet bulk Lorentz factor, which is determined by the jet dynamics. The catch-up radius is given by $r_{\rm c}=v_{\rm t}t_{\rm c}$. The jet choke condition is therefore $r_{\rm c}<r_{\rm ph}$.
During the jet propagation in ambient medium, two shocks are generated, the forward and reverse shocks (FS and RS). The shocks and the jet head can divide the system into four region: (1) unshocked ambient medium, (2) shocked ambient medium, (3) shocked shell material and (4) unshocked shell material.
In the observer frame, the radius of the deceleration $r_{\rm d}$ is a radius over which the accumulated circum-burst medium mass is $1/\Gamma_{0}$ of the ejecta isotropic mass (Zhang et al., 2003; Mészáros et al., 1998; Kobayashi, 2000), and the deceleration time $t_{\rm d}=r_{\rm d}/2\Gamma_{0}^{2}c$, where $\Gamma_{0}$ is the initial Lorentz factor of the jet and $E_{\rm iso}=L_{\rm j}T_{\rm j}/(1-\cos\theta_{\rm j}$). The Lorentz factors of region 2 and 3 are (Sari & Piran, 1995),
$\gamma_{2}=\gamma_{3}\approx\left\{\begin{aligned} &\Gamma_{0},&t_{\rm d}<T_{\rm j}~\text{(thin shell regime)}\\ &\Gamma_{0}^{1/2}f^{1/4}/\sqrt{2},&t_{\rm d}>T_{\rm j}~\text{(thick shell regime)}\end{aligned}\right$. (15)
where $f=n_{4}/n_{1}$ is the number density ratio between region 4 and 1, and $n_{4}=E_{\rm iso}/(4\pi m_{\rm p}c^{2}r^{2}\Gamma_{0}^{2}\Delta_{\rm j})$ and $n_{1}$ are the number density of the region 4 and 1, and $\Gamma_{\rm j}\sim\gamma_{2}$. The choke time will be $t_{\rm c}=t_{\rm d}$ for the thin shell regime, and $t_{\rm c}=T_{\rm j}$ for the thick shell regime.
@@ 3 Results
The jet break out from the isotropic wind have been studied by Teboul & Metzger (2023) and Lu et al. (2024). The ZEBRA envelope contains a low density polar funnels where jet can freely escape, and is most likely inclined. We need to check whether the precessing jet can successfully break out from such inclined ZEBRA envelope. The parameter determining the geometry ($\theta_{\rm LT}$, $\theta_{\rm env}$, $\theta_{\rm j}$, $\phi_{\rm obs}$), the jet dynamics ($P_{\rm LT}$, $\epsilon$, $\Gamma_{0}$), envelope structure ($M_{\bullet}$ and $M_{*}$) will come into consideration.
In the calculation, we assume the polar angle of observer is equal to the precessing angle of the jet unless otherwise specified, $\theta_{\rm obs}=\theta_{\rm LT}$. We assume that the initial Lorentz factor of jet is $\Gamma_{0}=10$. We assume that the envelope properties remain nearly unchanged during the jet precession (The precessing period is usually much smaller than the TDE time).
@@
Figure 3: The logarithm of the ratio between the reverse-shock crossing radius and the photosphere of the envelope for the jet, $\log(r_{\rm c}/r_{\rm ph}$). Successful breakout jet requires $\log(r_{\rm c}/r_{\rm ph})>0$, as shown by the red area. panel (a) is the fiducial case for $M_{\bullet}=10^{6}M_{\odot}$, t=100dayt=100~day, $\theta_{\rm j}=0.1$, $\epsilon=0.01$, $P_{\rm LT}=5~days$, $M_{*}=M_{\odot}$ and $\theta_{\rm LT}=\theta_{\rm LT}\in(0,\pi/2)$. In other panels, we only change one parameter for comparing with fiducial case; Panel (b), $\epsilon=0.05$, panel (c), $M_{\bullet}=10^{5}M_{\odot}$, panel (d), t=500dayt=500~day; panel (e), $\theta_{\rm j}=0.4$; panel (f), $P_{\rm LT}=10~days$; panel (g), $M_{*}=0.5M_{\odot}$; panel (h), $\theta_{\rm LT}<\theta_{\rm env}=0.5$; panel (i), $\theta_{\rm LT}<\theta_{\rm env}=1.5$. Besides, since larger $M_{\bullet}$ leads to a later onset of envelope evolution, we adopt the time of the peak $q$ to weaken the influence of the time.
@@ 4.3 Comparison with Observations
    Swift J1644+57: a jetted TDE with precessing.
    The X-ray temporal evolution of J1644 exhibits periodical dips (Bloom et al., 2011; Burrows et al., 2011; Levan et al., 2011). Within our framework, the early observed precession may be the case when the line of sight is just along the ZEBRA polar axis and the jet precess with $\theta_{\rm LT}\sim\theta_{\rm env}$. During the late evolutionary phase of the ZEBRA, the accretion rate declines to the Eddington limit, leading to the shut-off of relativistic jet (Coughlin & Begelman, 2014), which corresponds to the steep decaying light curve in the late-time emission. The absence of detectable optical emission from J1644 is attributed to significant dust extinction within its host galaxy, which effectively suppresses the intrinsic UV/optical continuum radiation (Burrows et al., 2011; Levan et al., 2011).
@@
    IGR J12580: a TDE with an off-beam relativistic jet.
    IGR J12580 initially observed by Integral in had X-ray band (Walter et al., 2011). Follow-up observations revealed that the X-ray light curve of J12580 maintained its peak luminosity for several tens of days before transitioning to $t^{-5/3}$ decay. While the overall evolutionary behavior resembles that of J1644, J12580 exhibits less variability in X-ray lightcurve (Burrows et al., 2011; Zauderer et al., 2013). Approximately one year after its X-ray discovery, the corresponding radio afterglow of J12580 was detected by the Karl G. Jansky Very Large Array (JVLA) (Irwin et al., 2015). Within our model framework, J12580 display a different situation with J1644, where the line of sight is away from the successful jet. The X-ray emission from the off-beam relativistic jet is strongly suppressed (Lei et al., 2016; Yuan et al., 2016). The radio emission from the off-beam jet become visible when the jet is decelerated to non-relativistic (Lei et al., 2016; Yuan et al., 2016).
@@
    AT2022cmc: a jetted TDE without precessing.
    Unlike J1644, no precession signatures have been observed in AT2022cmc. Observations show that its X-ray light curve rapidly transitions into a power-law decay phase during the early evolution (Andreoni et al., 2022). Within our theoretical framework, the absence of early jet precessing signatures arises from the alignment of jet, envelope polar axis, and SMBH spin axis. The line of sight is just along the spin axis, observing a bright and non-precessing jet emission.
@@
    ASASSN-14li: a TDE with the radio emission and fluctuation in UV/X-ray lightcurve.
    ASASSN-14li was initially discovered through optical emission, followed by subsequent detections across X-ray, UV, and radio emission (Miller et al., 2015; Holoien et al., 2016). The existence of a weak jet is supported by the correlation between radio and X-ray flares (Pasham & van Velzen, 2018). In our model, the jet may precess with a small precession angle in a ZEBRA envelope with small inclination, which make the jet more easy to break out. The breakout leaves behind a weak jet as observed in ASASSN-14li.

### v2

@@ 1 Introduction
A star would be disrupted by a supermassive black hole (SMBH) when it comes too close to the central SMBH. In such events, namely tidal disruption events (TDEs), the stellar debris are eventually accreted by the SMBH, producing bright multi-band emission (Hills, 1975; Rees, 1988). The initial stellar orbit is most likely misaligned with the equatorial plane of the spinning SMBH. The tilted disk formed after the TDE would undergo Lense-Thirring precssion around the SMBH spin axis (e.g., Papaloizou & Lin, 1995; Fragile et al., 2007; Stone & Loeb, 2012; Lei et al., 2013; Wang et al., 2014; Franchini et al., 2016; Li et al., 2025). The $\sim 15$ day quasi-periodic modulations found in the high-cadence X-ray monitoring observations of TDE candidate AT2020ocn, suggest the Lense-Thirring precession of a newly formed accretion disk (Pasham et al., 2024).
@@
For jetted TDEs, a question emerges: why would only a small fraction of TDEs successfully launch jets? One factor is the narrow beaming cone of the jets, and our line of sight is usually away from such jet cones. A recent study on IGR J12580+0134 revealed the first TDE with an off-axis relativistic jet (Lei et al., 2016; Yuan et al., 2016). Dai et al. (2018) proposed a unified model based on 3D general relativistic radiation magneto-hydrodynamis (GRRMHD) simulation that can explain both the optical and X-ray TDEs. The observed emission type mainly depends on the viewing angle of the observer with respect to the disk orientation. X-ray emissions will be observed from the optically thin funnel surrounded by winds when viewed face-on. For edge-on observations, only optical/UV emissions can be detected due to the heavily reprocessing of X-ray emissions in the chick wind and disk (Dai et al., 2018; Thomsen et al., 2022; Chen et al., 2025). However, even after the correction of relativistic beaming of jet and view angle effects, the fraction of jetted TDEs ($10^{-3}$ to $10^{-2}$ of all TDEs) is still much smaller than the fraction of active galactic nuclei with jets ($\sim 10$ per cent) (Teboul & Metzger, 2023; Lu et al., 2024). This remains an open question for jetted TDEs.
@@ 2.1 ZEBRA Envelope
where q=3/2nq=3/2-n, $\alpha=\frac{1-q(\gamma-1)}{\gamma-1}$, $\beta=\frac{\gamma-1}{1+\gamma-q(\gamma-1)}$, and $a=\frac{2-2q(\gamma-1)}{1+\gamma-q(\gamma-1)}$. $r_{0}\sim\chi r_{\rm s}$ is the inner radius of the envelope. Here, $r_{\rm s}=2GM_{\bullet}/c^{2}$ is the Schwarzschild radius, and $\chi$ is order of a few. $\rho_{0}$ is the density at this radius. The parameter $n$ is defined by Blandford and Begelman so that the accretion rate $\propto r^{n}$.
@@
Fig. 1 illustrates the time evolution of the $q$ and the global properties of the ZEBRA envelope. We consider the disruption of star with mass $M_{*}=M_{\odot}$ (solid lines) and $M_{*}=0.5M_{\odot}$ (dashed lines) by SMBH with mass $M_{\bullet}=10^{5}M_{\odot}$ (black lines) and $M_{\bullet}=10^{6}M_{\odot}$ (red lines). We show the evolution of $q$ in the panel (a). One can find that $M_{\bullet}$ and $M_{*}$ have significant effects for the whole evolution of $q$. In the panels (b)-(e), we also display the evolution of $T_{\rm env}$, $\rho_{0}$, $L_{\rm j}$ and $r_{\rm out}$ for different $M_{*}$ and $M_{\bullet}$. In addition, one should note that the accretion rate of SMBH would be lower than the Eddington accretion rate for $M_{\bullet}\sim 10^{7}M_{\odot}$(Coughlin & Begelman, 2014), and the ZEBRA envelope may not be formed. Therefore we do not consider such case in this work.
@@ 2.2 Precessing Jet
Figure 2: A schematic picture of the precession jet in an inclined ZEBRA envelope, the central orange ring represents the inner accretion disk within the ZEBRA envelope. The jet precess with angle $\theta_{\rm LT}$. The angle between the SMBH spin axis ($\Omega_{\bullet}$) and the envelope’s rotation axis ($\Omega_{\rm env}$), i.e., the envelope inclination, is $\theta_{\rm env}$. Left panel: $\theta_{\rm LT}=\theta_{\rm env}$. In the low density polar regions, i.e, “free zone”, the relativistic jet can escape freely when it precess to this direction. The jet is choked at other regions. Middle and right panels: $\theta_{\rm LT}<\theta_{\rm env}$. The jet will be choked in all directions if $\theta_{\rm LT}$ is large (middle panel), or break out of envelope if $\theta_{\rm LT}$ is small (right panel).
We consider an inner disk locating at the center of the envelope (Loeb & Ulmer, 1997). The envelop is much large in size, and we assume that only the inner disk undergoes significant Lense-Thirring precession.
Now, we introduce a precessing jet (with opening angle $\theta_{\rm j}$) launching from the center inner disk of the inclined ZEBRA envelope. The inner accretion disk undergoes Lense-Thirring precession with an precessing angle $\theta_{\rm LT}\leqslant\theta_{\rm env}$. We assume that the jet is normal to the inner disk and precess with the same period $P_{\rm LT}$. We show the schematic picture of our model in Fig. 2.
Due to the super-Eddington accretion, the inner disk is expected to be thick (Ulmer, 1997), and the surface density of the disk follows $\Sigma\propto r^{\rm-\zeta}$. The inner radius of the disk $r_{\rm d,in}$ is the innermost stable circular orbit (ISCO) and the outer radius $r_{\rm d,out}$ is the circularization radius of the TDE debris. $P_{\rm LT}$ depends on the SMBH mass $M_{\bullet}$ and spin $a_{\bullet}$, the mass of disrupted star $M_{*}$ and the boundaries of the precessing disk. According to the method of Franchini et al. (2016) and Stone & Loeb (2012),
$P_{\rm LT}=\frac{8\pi GM_{\bullet}(1+2)\zeta}{c^{3}(5-2\zeta)}\frac{\xi_{\rm out}^{5/2-\zeta}\xi_{\rm in}^{1/2+\zeta}\left[1-(\frac{\xi_{\rm in}}{\xi_{\rm out}})^{5/2-\zeta}\right]}{a_{\bullet}\left[1-(\frac{\xi_{\rm in}}{\xi_{\rm out}})^{1/2+\zeta}\right]},$
where $\xi_{\rm in}=r_{\rm d,in}/r_{\rm s}$ and $\xi_{\rm out}=r_{\rm d,out}/r_{\rm s}$. We calculate the $P_{\rm LT}$ evolution with SMBH spin $a_{\bullet}$ for $_{\bullet}\sim 10^{6}M_{\odot}, MM$$_{*}\sim M_{\odot}$ and $\zeta=0.6$, as shown in Fig.3. One can get $P_{\rm LT}\sim 40$~days for $a_{\bullet}\sim 0.1$, $P_{\rm LT}\sim 6$~days for $a_{\bullet}\sim 0.5$ and $P_{\rm LT}\sim 1$~day for $a_{\bullet}\sim 1$.
Figure 3: The precessing period $P_{\rm LT}$ as a function of SMBH spin $a_{\bullet}$ for $M_{\bullet}\sim 10^{6}M_{\odot}$ and $M_{*}\sim M_{\odot}$.
@@
The jet duration is $T_{\rm j}=\xi_{\rm duty}P_{\rm LT}$ and the width of the shell is $\Delta_{\rm j}\approx T_{\rm j}c$. The jet would be choked if its tail catches up to the head insider the ZEBRA envelope photosphere $r_{\rm ph}$ (defined by $\int_{r_{\rm ph}}^{r_{\rm out}}\kappa\rho dr=1$). Assuming the catch-up time $t_{\rm c}$ in the lab frame and the catch-up radius $r_{\rm c}$, one finds $(v_{\rm t}-v_{\rm h})t_{\rm c}=\Delta_{\rm j}$, where $v_{\rm t}\approx c$ and $v_{\rm h}\approx c\sqrt{1-1/\Gamma_{\rm j}^{2}}$ are the tail and head velocities of the jet, and $\Delta_{\rm j}\approx cT_{\rm j}$ is the length of the jet. The jet catch-up time is $t_{\rm c}=\Delta_{\rm j}/(v_{\rm t}-v_{\rm h})=T_{\rm j}/(\beta_{\rm t}-\beta_{\rm h})\approx T_{\rm j}/(1-\beta_{\rm h})\approx 2\Gamma_{\rm j}^{2}T_{\rm j}$, and $\Gamma_{\rm j}$ is the jet bulk Lorentz factor which is determined by the jet dynamics, $\beta_{\rm t}=v_{\rm t}/c$ and $\beta_{\rm h}=v_{\rm h}/c$. The catch-up radius is given by $r_{\rm c}=v_{\rm t}t_{\rm c}$. The jet choke condition is therefore $r_{\rm c}<r_{\rm ph}$.
During the jet propagation in ambient medium, e.g., the ZEBRA envelope medium, two shocks are generated, the forward and reverse shocks (FS and RS). The shocks and the jet head can divide the system into four region: (1) unshocked ZEBRA medium, (2) shocked ZEBRA medium, (3) shocked jet shell material and (4) unshocked jet shell material.
In the observer frame, the radius of the deceleration $r_{\rm dec}$ is a radius over which the accumulated ZEBRA medium mass is $1/\Gamma_{0}$ of the jet isotropic mass (Zhang et al., 2003; Mészáros et al., 1998; Kobayashi, 2000), which can be defined as $\int_{r_{0}}^{r_{\rm dec}}4\pi r^{2}\Gamma_{0}\rho dr=E_{\rm iso}/\Gamma_{0}c^{2}$, and the deceleration time $t_{\rm dec}=r_{\rm dec}/2\Gamma_{0}^{2}c$, where $\Gamma_{0}$ is the initial Lorentz factor of the jet and $E_{\rm iso}=L_{\rm j}T_{\rm j}/(1-\cos\theta_{\rm j}$). The Lorentz factors of region 2 and 3 are (Sari & Piran, 1995),
$\gamma_{2}=\gamma_{3}\approx\left\{\begin{aligned} &\Gamma_{0},&t_{\rm dec}<T_{\rm j}~\text{(thin shell regime)}\\ &\Gamma_{0}^{1/2}f^{1/4}/\sqrt{2},&t_{\rm dec}>T_{\rm j}~\text{(thick shell regime)}\end{aligned}\right$. (16)
where $f=n_{4}/n_{1}$ is the number density ratio between region 4 and 1, and $n_{4}=E_{\rm iso}/(4\pi m_{\rm p}c^{2}r^{2}\Gamma_{0}^{2}\Delta_{\rm j})$ and $n_{1}$ are the number density of the region 4 and 1, and $\Gamma_{\rm j}\sim\gamma_{2}$. The choke time will be $t_{\rm c}=t_{\rm dec}$ for the thin shell regime, and $t_{\rm c}=T_{\rm j}$ for the thick shell regime.
@@ 3 Results
The jet break out from the isotropic wind have been studied by Teboul & Metzger (2023) and Lu et al. (2024). We show their results in the Fig. 4 using red (for Teboul & Metzger (2023)) and purple (for Lu et al. (2024)) dashed lines, respectively. The ZEBRA envelope contains a low density polar funnels where jet can freely escape, and is most likely inclined. We need to check whether the precessing jet can successfully break out from such inclined ZEBRA envelope. The parameter determining the geometry ($\theta_{\rm LT}$, $\theta_{\rm env}$, $\theta_{\rm j}$, $\phi_{\rm obs}$), the jet dynamics ($P_{\rm LT}$, $\epsilon$, $\Gamma_{0}$), envelope structure ($M_{\bullet}$ and $M_{*}$) will come into consideration.
In the calculation, we assume the polar angle of observer is equal to the precessing angle of the jet unless otherwise specified, $\theta_{\rm obs}=\theta_{\rm LT}$. We assume that the initial Lorentz factor of jet is $\Gamma_{0}=10$ [^1] One should note that Coughlin & Begelman (2014) adopted the radiation powered jet. However, the numerical simulations by Sądowski & Narayan (2015) and Sądowski et al. (2015) found that the radiation alone cannot produce the high Lorentz factor of jet. The general relativistic radiation MHD simulations of accretion onto a SMBH during a TDE shows that a strong poloidal magnetic field and rapid BH spin are needed to produce the jet power in J1644 (Curd & Narayan, 2019). Therefore, we assume that the jet is produced by the influence of rapid SMBH spin and strong magnetic field, and adopt $\Gamma_{0}=10$ as a fiducial value in the calculation.. We assume that the envelope properties remain nearly unchanged during the jet precession (The precessing period is usually much smaller than the duration of TDE).
@@
Figure 4: The logarithm of the ratio between the reverse-shock crossing radius and the photosphere of the envelope for the jet, $\log(r_{\rm c}/r_{\rm ph}$). Successful breakout jet requires $\log(r_{\rm c}/r_{\rm ph})>0$, as shown by the red area. panel (a) is the fiducial case for $M_{\bullet}=10^{6}M_{\odot}$, $t=100~\rm day$, $\theta_{\rm j}=0.1$, $\epsilon=0.01$, $P_{\rm LT}=5~\rm day$, $M_{*}=M_{\odot}$ and $\theta_{\rm LT}=\theta_{\rm env}\in(0,\pi/2)$. In other panels, we only change one parameter for comparing with fiducial case; Panel (b), $\epsilon=0.05$, panel (c), $M_{\bullet}=10^{5}M_{\odot}$, panel (d), $t=500~\rm day$; panel (e), $\theta_{\rm j}=0.4$; panel (f), $P_{\rm LT}=10~\rm day$; panel (g), $M_{*}=0.5M_{\odot}$; panel (h), $\theta_{\rm LT}<\theta_{\rm env}=0.5$; panel (i), $\theta_{\rm LT}<\theta_{\rm env}=1.5$. Besides, since larger $M_{\bullet}$ leads to a later onset of envelope evolution, we adopt the time of the peak $q$ to weaken the influence of the time. The red and purple dashed lines are the results of Teboul & Metzger (2023) and Lu et al. (2024) for $\epsilon=0.01$ and $0.05$, respectively.
@@ 4.3 Comparison with Observations
In the following calculations, we assume $\phi_{\rm obs}\sim 0$ for simplicity. Besides, the structure of ZEBRA envelope that we adopt is that in the fiducial case (see panel a in Fig.4), i,e, $M_{\bullet}=10^{6}M_{\odot}$, $t=100~\rm day$ and $M_{*}=M_{\odot}$. We also take $\epsilon=0.01$ as in the fiducial case.
@@
    Swift J1644+57 (J1644): a jetted TDE with precessing.
    The X-ray temporal evolution of J1644 exhibits periodical dips (Bloom et al., 2011; Burrows et al., 2011; Levan et al., 2011). We adopt $P_{\rm LT}\sim 2.7$~day for J1644 (Lei et al., 2013). According to the Fig.3, one can estimate the spin of SMBH $a_{\bullet}\sim 0.75$. For jets with significant precession, the observed flux primarily depends on two factors: one is $\psi$, which is the angle between the line of sight and the jet axis, and the other is the jet Lorentz factors when the jet breaks out from the ZEBRA envelope. The evolution of $\psi$ leads to the on-axis and off-axis of the jet, i.e., the duty cycle $\xi_{\rm duty}$. For J1644, the duty cycle is roughly $\xi_{\rm duty}\sim 0.4$. Since we observed strong emission and high Lorentz factor in J1644 jet, we assume the jet escape along the ZEBRA polar ($\theta_{\rm LT}\sim\theta_{\rm env}$). Besides, we assume the opening angle of the jet $\theta_{\rm j}\sim 6^{\circ}$ (Lei et al., 2013). Another parameter is the peak-to-dip flux ratio $\lambda=F(\psi=\psi_{\rm max},t)/F(\psi=0,Dt)$, where $\psi_{\rm max}$ is the maximum value of $\psi$. The observed flux is given by $F(\psi,t)=D^{4}F(\psi=0,Dt)$, where $D=\Gamma_{\rm on}(1-\beta_{\rm on})/\Gamma_{\rm off}(1-\beta_{\rm off}\cos\psi)$ is the ratio of the off-beam Doppler factor to the on-beam Doppler factor, $\Gamma_{\rm on}=1/\sqrt{1-\beta_{\rm on}^{2}}$ ($\Gamma_{\rm off}=1/\sqrt{1-\beta_{\rm off}^{2}}$) is the Lorentz factor for $\psi=0$ ($\psi>0$). In the ZEBRA envelope, the density distribution is not isotropic. When the jet precesses to different orientations, it sweeps through different density medium, resulting in different Lorentz factors when the jet breaks out. We adopt the method described in Huang et al. (1999) to calculate the evolution of the jet Lorentz factor. We use the empirical relationship between $\Gamma$ and $L_{\rm j,iso}$ to derive the jet initial Lorentz factor $\Gamma_{0}$ (Wu et al., 2016), where $L_{\rm j,iso}=L_{\rm j}/(1-\cos\theta_{\rm j})=L_{\rm X,iso}/\epsilon_{\rm X}$ is the isotropic luminosity of the jet, $L_{\rm X,iso}$ is the isotropic radiation luminosity in X-ray and $\epsilon_{\rm X}\sim 0.1$ is the efficiency of jet luminosity to X-ray radiation. For J1644, $L_{\rm X,iso}\sim 10^{47}~\rm erg~s^{-1}$ and $\Gamma_{0}\sim 30$. We also use $\lambda\sim 0.3$ (Lei et al., 2013). The line of sight $\theta_{\rm obs}$ and precession angle $\theta_{\rm LT}$ should satisfy the constraints both $\xi_{\rm duty}$ and $\lambda$, see Fig.5. One can find two different solutions, e.g., ($\theta_{\rm obs}=1.19^{\circ}$, $\theta_{\rm LT}=6.26^{\circ}$) and ($\theta_{\rm obs}=6.26^{\circ}$, $\theta_{\rm LT}=1.19^{\circ}$). J1644’s jet is initially precessing but ceased after $\sim 30$ days due to the alignment with the SMBH spin (Lei et al., 2013; Teboul & Metzger, 2023). The jet emission is observed even after the alignment, we then expect a small $\theta_{\rm obs}$ (otherwise, the line of sight would be outside the jet cone after the alignment). Therefore, for J1644, one can get $\theta_{\rm obs}=1.19^{\circ}$ and $\theta_{\rm LT}=6.26^{\circ}$. During the late evolutionary phase of the ZEBRA, the accretion rate declines to the Eddington limit, leading to the shut-off of relativistic jet (Coughlin & Begelman, 2014), which corresponds to the steep decaying light curve in the late-time emission. The absence of detectable optical emission from J1644 is attributed to significant dust extinction within its host galaxy, which effectively suppresses the intrinsic UV/optical continuum radiation (Burrows et al., 2011; Levan et al., 2011).
@@
    AT2022cmc: a jetted TDE without precessing.
    Unlike J1644, no precession signatures have been observed in AT2022cmc. Observations show that its X-ray light curve rapidly transitions into a power-law decay phase during the early evolution (Andreoni et al., 2022). Within our theoretical framework, the absence of early jet precessing signatures arises from the nearly alignment of jet, envelope polar axis, and SMBH spin axis. Since we observe strong jet emission but without significant precession in this case, we assume the jet escapes along the ZEBRA polar axis ($\theta_{\rm LT}\sim\theta_{\rm env}$), but expect a small jet precesssing angle $\theta_{\rm LT}$ and observer angle $\theta_{\rm obs}$. We adopt $\theta_{\rm j}\sim 6^{\circ}$, $L_{\rm x,iso}\sim 5\times 10^{46}~\rm erg~s^{-1}$, $\xi_{\rm duty}\sim 0.75$ and $\lambda\sim 0.85$ to estimate the values of $\theta_{\rm LT}$ and $\theta_{\rm obs}$ with the similar method for the J1644, as shown in the panel (b) of Fig.5. We also use the $\Gamma-L_{\rm X,iso}$ empirical relation (Wu et al., 2016) to get $\Gamma_{0}\sim 25$ for AT2022cmc. We find $\theta_{\rm obs}\sim 3.21^{\circ}$ and $\theta_{\rm LT}\sim 3.37^{\circ}$.
@@
    IGR J12580 (J12580): a TDE with an off-beam relativistic jet.
    IGR J12580 initially observed by Integral in had X-ray band (Walter et al., 2011). Follow-up observations revealed that the X-ray light curve of J12580 maintained its peak luminosity for several tens of days before transitioning to $t^{-5/3}$ decay. While the overall evolutionary behavior resembles that of J1644, J12580 exhibits less variability in X-ray lightcurve (Burrows et al., 2011; Zauderer et al., 2013). Approximately one year after its X-ray discovery, the corresponding radio afterglow of J12580 was detected by the Karl G. Jansky Very Large Array (JVLA) (Irwin et al., 2015). We expect a successful jet from J12580, which escapes from the polar axis of ZEBRA envelope. We therefore assume $\theta_{\rm LT}\sim\theta_{\rm env}$. The X-ray emission from the off-beam relativistic jet is strongly suppressed (Lei et al., 2016; Yuan et al., 2016), and the radiation luminosity is $\sim 1/7200$ of J1644. In addition, we assume $\theta_{\rm j}\sim 8^{\circ}$ and $\Gamma_{0}\sim 5$ to estimate $\theta_{\rm obs}$ and $\theta_{\rm LT}$ for $\lambda=1/7200$. The result is shown in the panel (c) of Fig.5. Since only one constraint is applied, all points along the blue solid line are potential solutions. Since no significant precession observed, we expect the jet nearly aligns with the SMBH spin ($\theta_{\rm LT}\sim 0$), and the solution with large $\theta_{\rm obs}$ ($\sim 40^{\circ}$) will be consistent with Lei et al. (2016). The radio emission from the off-beam jet become visible when the jet is decelerated to non-relativistic (Lei et al., 2016; Yuan et al., 2016).
@@
    ASASSN-14li: a TDE with the UV/radio emission and thermal X-ray component.
    ASASSN-14li was initially discovered through optical emission, followed by subsequent detections across X-ray, UV, and radio emission, and X-ray emission is dominated by thermal emission(Miller et al., 2015; Holoien et al., 2016). The existence of a weak jet is supported by the correlation between radio and X-ray flares (Pasham & van Velzen, 2018). In our model, the jet is weak (Pasham & van Velzen, 2018). Due to no significant preceesing signal in the UV band emission, we assume the jet is aligned with the SMBH spin and the line of sight is also along the SMBH spin.